# Fairshare Data Pricing for Large Language Models

### v1

@@ Appendix D Proofs
Lemma 3 Suppose we have a LLM with parameters $\theta$. We perform a gradient descent step with training sample $d$ with learning rate $\eta$ such that $\hat{\theta}=\theta-\eta\nabla\mathcal{L}(d;\theta)$. Then,
@@
We show that, for every $\boldsymbol{x}\in\mathcal{X}_{k,j-1,\text{frac}}^{1}$, $\alpha_{j}^{\boldsymbol{x}}$ gives the largest revenue of $\boldsymbol{x}^{\text{\text{new}}}$ for ${S}_{j}$. Recall that in the main text, we need to increase $\alpha_{j}$ from zero until we find the largest $\alpha_{j}$ such that either of:
@@
    $g_{k,j,\text{frac}}(\boldsymbol{x}^{\text{\text{new}}})>g_{k,j,\text{frac}}(\tilde{\boldsymbol{x}}^{k,j-1,\text{frac}})$,
@@
    $f_{j}(\begin{bmatrix}\boldsymbol{\alpha}^{T}\;\alpha_{j}\end{bmatrix},\boldsymbol{x}^{\text{\text{new}}})=\overline{\alpha}_{k}$.
@@
$\sup_{\alpha_{j}\in[0,1)}\left\{\alpha_{j}:f_{j}(\begin{bmatrix}\boldsymbol{\alpha}^{T}\;\alpha_{j}\end{bmatrix},\boldsymbol{x}^{\text{\text{new}}})<1-(1-f_{j}(\boldsymbol{\alpha},\tilde{\boldsymbol{x}}^{k,j-1,\text{frac}}))\frac{u_{k}(\tilde{\boldsymbol{x}}^{k,j-1,\text{frac}})}{u_{k}(\boldsymbol{x}^{\text{\text{new}}})}\right\}$ (42)
Then we see that the revenue that for each $\boldsymbol{x}\in\mathcal{X}_{k,j-1,\text{frac}}^{1}$, seller ${S}_{j}$ can make from buyer $B_{k}$ is
$f_{j}(\begin{bmatrix}\boldsymbol{\alpha}^{T}\;\alpha_{j}\end{bmatrix},\boldsymbol{x}^{\text{\text{new}}})u_{k}(\boldsymbol{x}_{j}(\boldsymbol{\alpha})),$ (43)
where $f_{j}(\begin{bmatrix}\boldsymbol{\alpha}^{T}\;\alpha_{j}\end{bmatrix},\boldsymbol{x}^{\text{\text{new}}})$ is a non-decreasing function over $\alpha_{j}$ while other terms stays fixed. It indicates that $\alpha_{j}^{\boldsymbol{x}}$ is the largest $\alpha_{j}$ that the seller ${S}_{j}$ could set for buyer $B_{k}$ to purchase $S_{j}$. Therefore, the optimal rate $\alpha_{j}^{*}$ is one of the rates $\cup_{k=1}^{M}\cup_{\boldsymbol{x}\in\mathcal{X}_{k,j-1,\text{frac}}^{1}}\{\alpha_{j}^{\boldsymbol{x}}\}$. ∎
@@ E.1 Data Valuation Experiments
Model Training: After obtaining purchasing decisions for all data samples, the buyers train their models using the purchased data. In order to conduct a fair comparison across buyers, we sample a set number of data from the buyers’ purchases (shown in Figure 3). We train each model (i.e., buyer) on these samples separately using LoRA (Hu et al., 2021) for 3 epochs, with a learning rate of 2e-7 and batch size 32.
@@
Market/Pricing Setup: We reserve $1\%$ of the samples from each dataset’s training split to represent the existing data in their respective markets. Each data sample was randomly priced between $(0,1]$. Next, for each remaining data sample in the training set, we determine whether each buyer will purchase the data sample at potential price points $[0.5,0.625,0.75,0.875,1.0]$ by solving Equation 2. The seller then sets their prices according to Equation 5. We price data separately for each data valuation method. This assesses the method’s ability to discern whether a new data sample is worth purchasing for each buyer given the existing market data, as noted in our analysis in Section 3.2.
@@ E.2 Data Pricing Experiments
Experiment Setups: We simulate two buyer budgets at each time step $t$. The first buyer (high budget) has a budget uniformly randomly generated between $95\%$ and $100\%$ of the total utilities of all $10$ datasets listed in the market. The second buyer (low budget) has a budget uniformly randomly generated between $90\%$ and $95\%$ of the total utilities of all $10$ datasets listed in the market.
Appendix F Additional Tables and Figures
@@ F.1 Algorithms
1:  Inputs: prices $\boldsymbol{p}$, optimal solution $\tilde{\boldsymbol{x}}^{k,j-1}$, feasible solutions $\mathcal{X}_{k,j-1}$, price $p_{j}$.
2:  Output: $\mathbbm{1}_{\{B_{k},D_{j},p_{j}\}}$.
3:  Initialize $\mathbbm{1}_{\{B_{k},D_{j},p_{j}\}}\rightarrow 0$.
4:  for $\boldsymbol{x}\in\mathcal{X}_{k,j-1}$ do
5:     if $g_{k,j}(\boldsymbol{x}^{\text{new}})>g_{k,j-1}(\tilde{\boldsymbol{x}}^{k,j-1})$ and $\boldsymbol{x}^{T}\boldsymbol{p}+p_{j}\leq b_{k}$ then
6:        $\mathbbm{1}_{\{B_{k},D_{j},p_{j}\}}\leftarrow 1$
7:     end if
8:  end for
9:  Return $\mathbbm{1}_{\{B_{k},D_{j},p_{j}\}}$
1:  Inputs: Buyers $\{B_{k}\}_{k=1}^{M}$ and sellers $\{S_{j}\}_{j=1}^{N}$.
2:  Initialization: Buyers $\{B_{k}\}_{k=1}^{M}$ enter the market.
3:  for $j\in[N]$ do
4:     $S_{j}$ enters the market with potential prices $\mathcal{P}_{j,M}$ for dataset $D_{j}$
5:     for $p_{j}\in\mathcal{P}_{j,M}$ do
6:        for $k\in[M]$ do
7:           $B_{k}$ solves for $\tilde{\boldsymbol{x}}^{k,j-1}=\operatorname*{arg\,max}_{\boldsymbol{x}\in\mathcal{X}_{k,j}}\;g_{k,j}(\boldsymbol{x})$ to determine if they will purchase $D_{j}$ at potential price $p_{j}$. (Eqn. 2)
8:        end for
9:        Seller calculates net profit $r(p_{j})$ if they sold potential price $p_{j}$. (Eqn. 4)
10:     end for
11:     Seller solves for $p^{*}_{j}=\operatorname*{arg\,max}_{{p}_{j}\in\mathcal{P}_{j,M}}\;r(p_{j})$, and sets $p^{*}_{j}$ as the price for $D_{j}$ (Eqn. 5), which is fixed for subsequent rounds.
12:  end for
@@
F.2 Datasets

### v2

@@ Appendix D Lemmas and Proofs
Lemma 6 (Participation Loss Is Lower-Bounded).
With Assumption 1, let $P:=\lim_{T\rightarrow\infty}P_{T}$ and $S:=\sum_{i=0}^{\infty}(1-\pi(p_{i},p^{*}_{i}))$. Then for the class of all exploitative pricing strategies where $p_{t}\leq p_{t}^{*},\forall t$, we have (1) $P_{T}$ is strictly decreasing and (2) The limit of reduced participation is lower bounded, i.e., $01-P\geq 1-e^{-S}>$ 0.
Proof.
The first part is trivial to see as we assume that $\pi(p_{i},p^{*}_{i})<1$ for all $p_{i}<p^{*}_{i}$ from Assumption 1.
For the second part, we first denote $P:=\lim_{t\rightarrow\infty}P_{t}$. Then we take the following transformation:
$\log P=\log\left(\prod_{i=0}^{\infty}\pi(p_{i},p^{*}_{i})\right)=\sum_{i=0}^{\infty}\log\left(\pi(p_{i},p^{*}_{i})\right).$ (31)
Since we have $10<\pi(p_{i},p^{*}_{i})<$ 1, then $\log\left(\pi(p_{i},p^{*}_{i})\right)\leq\pi(p_{i},p^{*}_{i})-1$, indicating that
$\log P=\sum_{i=0}^{\infty}\log\left(\pi(p_{i},p^{*}_{i})\right)\leq\sum_{i=0}^{\infty}\left(\pi(p_{i},p^{*}_{i})-1\right)=-S,$ (32)
which gives $P\leq e^{-S}$ by exponentiating both sides. ∎
Lemma 2. Seller $S_{j}$’s optimal price for $D_{j}$ is characterized as one of buyers’ MWP:
$p_{j}^{*}\in\cup_{k=1}^{M}\text{MWP}_{k}.$ (33)
Proof.
Recall that without dataset $D_{j}$, each buyer $B_{k}$ has already solved $\max_{\boldsymbol{x}\in\mathcal{X}_{k,N-1}}g_{k,N-1}(\boldsymbol{x})$ according to our market definition in Section 3, where $\mathcal{X}_{k,N-1}$ is the set of all feasible purchase decisions. Next, after seller $S_{j}$ (with $D_{j}$) has arrived on the market, we analyze the conditions in which $B_{k}$ will purchase $D_{j}$ at a potential price $p_{j}$. For each feasible purchase decision (i.e., a collection of datasets), represented by $\boldsymbol{x}\in\mathcal{X}_{k,N-1}$, let $\boldsymbol{x}+e_{j}$ denote its union with $D_{j}$, where $e_{j}$ is the unit vector indicating $D_{j}$ is selected. For buyer $B_{k}$ to change their previous decision to purchase $D_{j}$, there are two requirements that need to be satisfied. First, we must have:
$g_{k,N}(\boldsymbol{x}+e_{j})>g_{k,N-1}(\tilde{\boldsymbol{x}}^{k,N-1}).$ (34)
That is, the net utility $g_{k,N}(\boldsymbol{x}+e_{j})$ of purchasing decisions $\boldsymbol{x}+e_{j}$, must be larger than the net utility $g_{k,N-1}(\tilde{\boldsymbol{x}}^{k,N-1})$ of a previous optimal purchasing decision $\tilde{\boldsymbol{x}}^{k,N-1}$. It is also noted that $g_{k,N}(\tilde{\boldsymbol{x}}^{k,N-1})=g_{k,N-1}(\tilde{\boldsymbol{x}}^{k,N-1})$. Second, for buyer $B_{k}$ to purchase $\boldsymbol{x}+e_{j}$ at price $p_{j}$, we must fulfill the budget constraint:
$p_{j}\leq b_{k}-\left(\tilde{\boldsymbol{x}}^{k,N-1}\right)^{T}\boldsymbol{p}=\Delta b_{k}(\tilde{\boldsymbol{x}}^{k,N-1}),$ (35)
which ensures that purchasing $D_{j}$ does not exceed the buyer’s budget $b_{k}$. If both requirements are satisfied, then the buyer $B_{k}$ will change their previous purchasing decision in order to purchase $D_{j}$ under the price $p_{j}$. This procedure is presented in detail in Algorithm 1 in Section F.1.
Next, given the conditions for the buyer $B_{k}$ to purchase $D_{j}$, the seller must solve $\max_{{p}_{j}\in\mathcal{P}_{j,M}}r(p_{j})$ to find the optimal price $p_{j}^{*}$. First, we consider an edge case where the price of dataset $D_{j}$ is set as $p_{j}=0$. For a buyer $B_{k}$, we denote $\mathcal{X}_{k,N}^{1}$ as the set of all purchasing decisions where including $D_{j}$ in the purchase improves the buyer’s previous net utility $g_{k,N}(\tilde{\boldsymbol{x}}^{k,N-1})$. That is, for every $\boldsymbol{x}+e_{j}\in\mathcal{X}_{k,N}^{1}$, we have $g_{k,N}(\boldsymbol{x}+e_{j})>g_{k,N}(\tilde{\boldsymbol{x}}^{k,N-1})$. If $\mathcal{X}_{k,N}^{1}$ is empty, then $B_{k}$ will not purchase $D_{j}$ at any price, since $D_{j}$ cannot bring positive improved net utility to $B_{k}$. Then, when $p_{j}$ gradually increases and exceeds $\max_{\boldsymbol{x}\in\mathcal{X}_{k,N-1}}\{\min\{\Delta u_{k}(\boldsymbol{x}+e_{j}),\Delta b_{k}(\tilde{\boldsymbol{x}}^{k,N-1})\}\}$, then $B_{k}$ will decide not to purchase $D_{j}$, causing the value of $\sum_{k=1}^{M}\tilde{\boldsymbol{x}}^{k,N}_{j}(\boldsymbol{p})$ to drop by one. Since the profit function $r(p_{j})$ is a piecewise linear function, the its optimal point must be one of its breakpoints. ∎
Lemma 3 The optimal price for the seller $S$ under our framework is
$p^{*}_{t}=\min\{u_{t},b_{t}\},\forall t.$ (36)
With assumptions 1, 1.1 and 2, $p^{*}_{t}$ gives the buyer the maximum cumulative net utility over infinite horizon.
Proof.
In a single buyer and seller setting, we could trivially see that the optimal price for the seller $p_{t}^{*}=\min\{u_{t},b_{t}\}$: if $p_{t}>p_{t}^{*}$, then the buyer would not purchase this dataset since the net utility would be negative; if $p_{t}<p_{t}^{*}$, then the seller’s profit is not maximized.
Further, we refer to the proof of Lemma 1 for the second part. ∎
Lemma 4. Suppose we have a LLM with parameters $\theta$. We perform a gradient descent step with training sample $d$ with learning rate $\eta$ such that $\hat{\theta}=\theta-\eta\nabla\mathcal{L}(d;\theta)$. Then,
@@
Lemma 5 Define $\alpha_{j}^{\boldsymbol{x}}$ as
$\min$$\left\{\sup_{\alpha_{j}\in[0,1)}\left\{\alpha_{j}:f_{j}(\begin{bmatrix}\boldsymbol{\alpha}^{T}\;\alpha_{j}\end{bmatrix},\boldsymbol{x}^{\text{new}})<1-(1-f_{j}(\boldsymbol{\alpha},\tilde{\boldsymbol{x}}^{k,N-1,\text{frac}}))\frac{u_{k}(\tilde{\boldsymbol{x}}^{k,N-1,\text{frac}})}{u_{k}(\boldsymbol{x}^{\text{new}})}\right\},\right.$
For every $\boldsymbol{x}\in\mathcal{X}_{k,N-1,\text{frac}}^{1}$ and all $k\in[M]$, we obtain $\alpha_{j}^{\boldsymbol{x}}$ and their union $\cup_{k=1}^{M}\cup_{\boldsymbol{x}\in\mathcal{X}_{k,N-1,\text{frac}}^{1}}\{\alpha_{j}^{\boldsymbol{x}}\}$. Then we have $\alpha_{j}^{*}\in\cup_{k=1}^{M}\cup_{\boldsymbol{x}\in\mathcal{X}_{k,N-1,\text{frac}}^{1}}\{\alpha_{j}^{\boldsymbol{x}}\}$.
@@
We show that, for every $\boldsymbol{x}\in\mathcal{X}_{k,N-1,\text{frac}}^{1}$, $\alpha_{j}^{\boldsymbol{x}}$ gives the largest revenue of $\boldsymbol{x}+e_{j}$ for ${S}_{j}$ (note that $\boldsymbol{x}+e_{j}=\boldsymbol{x}^{\text{new}}$). Recall that in the main text, we need to increase $\alpha_{j}$ from zero until we find the largest $\alpha_{j}$ such that either of:
@@
    $g_{k,N,\text{frac}}(\boldsymbol{x}+e_{j})>g_{k,N,\text{frac}}(\tilde{\boldsymbol{x}}^{k,N-1,\text{frac}})$,
@@
    $f_{j}(\begin{bmatrix}\boldsymbol{\alpha}^{T}\;\alpha_{j}\end{bmatrix},\boldsymbol{x}+e_{j})=\overline{\alpha}_{k}$.
@@
$\sup_{\alpha_{j}\in[0,1)}\left\{\alpha_{j}:f_{j}(\begin{bmatrix}\boldsymbol{\alpha}^{T}\;\alpha_{j}\end{bmatrix},\boldsymbol{x}+e_{j})<1-(1-f_{j}(\boldsymbol{\alpha},\tilde{\boldsymbol{x}}^{k,N-1,\text{frac}}))\frac{u_{k}(\tilde{\boldsymbol{x}}^{k,N-1,\text{frac}})}{u_{k}(\boldsymbol{x}+e_{j})}\right\}$ (41)
Then we see that the revenue that for each $\boldsymbol{x}\in\mathcal{X}_{k,N-1,\text{frac}}^{1}$, seller ${S}_{j}$ can make from buyer $B_{k}$ is
$f_{j}(\begin{bmatrix}\boldsymbol{\alpha}^{T}\;\alpha_{j}\end{bmatrix},\boldsymbol{x}+e_{j})u_{k}(\boldsymbol{x}_{j}(\boldsymbol{\alpha})),$ (42)
where $f_{j}(\begin{bmatrix}\boldsymbol{\alpha}^{T}\;\alpha_{j}\end{bmatrix},\boldsymbol{x}+e_{j})$ is a non-decreasing function over $\alpha_{j}$ while other terms stays fixed. It indicates that $\alpha_{j}^{\boldsymbol{x}}$ is the largest $\alpha_{j}$ that the seller ${S}_{j}$ could set for buyer $B_{k}$ to purchase $S_{j}$. Therefore, the optimal rate $\alpha_{j}^{*}$ is one of the rates $\cup_{k=1}^{M}\cup_{\boldsymbol{x}\in\mathcal{X}_{k,N-1,\text{frac}}^{1}}\{\alpha_{j}^{\boldsymbol{x}}\}$. ∎
Lemma 7 (The Trade-Off Threshold Is Increasing as $\delta$ Decreases).
The threshold time period where the fairshare pricing obtains higher cumulative utility than any class of exploitative pricing is:
$t^{*}:=\sup_{p_{t}<p_{t}^{*},\forall t}\left\{T\in\mathbb{N}:\mathbb{E}\left[\sum_{t=0}^{T}\delta^{t}\left(\left(u_{t}-p^{*}_{t}\right)-\prod_{i=0}^{t-1}\pi(p_{i},p^{*}_{i})\left(u_{i}-p_{i}\right)\right)\right]\leq 0\right\}.$ (43)
And $t^{*}$ is increasing as $\delta$ increases.
Proof.
For any given class of exploitative pricing strategy, when $\delta$ increases, the part: $\sum_{t=0}^{T}\delta^{t}\left(\left(u_{t}-p^{*}_{t}\right)-\prod_{i=0}^{t-1}\pi(p_{i},p^{*}_{i})\left(u_{i}-p_{i}\right)\right)$ increases. Therefore, for all class of exploitative pricing strategy, i.e., $p_{t}<p_{t}^{*},\forall t$, then $t^{*}$ also decreases. ∎
@@ E.1 Data Valuation Experiments
For each dataset, we randomly sample 200 demonstrations from the validation set to form a representative dataset  [^4] Note: For PIQA we take 200 samples from the training set since the validation set is commonly reserved for testing.. Each data sample in the market is then scored based on its similarity to this representative set.
Model Training: After obtaining purchasing decisions for all data samples, the buyers train their models using the purchased data. In order to conduct a fair comparison across buyers, we sample a set number of data from the buyers’ purchases (shown in Figure 6). We train each model (i.e., buyer) on these samples separately using LoRA [82] for 3 epochs, with a learning rate of 2e-5 and batch size 32. All models are trained on A6000 GPUs on single GPU settings and take less than 1 hour.
@@
Market/Pricing Setup: We reserve $1\%$ of the samples from each dataset’s training split to represent the existing data in their respective markets. Each data sample was randomly priced between $(0,1]$. Next, for each remaining data sample in the training set, we determine whether each buyer will purchase the data sample at potential price points $[0.5,0.625,0.75,0.875,1.0]$ by solving Equation 2. The seller then sets their prices according to Equation 4. We price data separately for each data valuation method. This assesses the method’s ability to discern whether a new data sample is worth purchasing for each buyer given the existing market data, as noted in our analysis in Section 3.
@@ E.2 Data Pricing Experiments
Experiment Setups: We simulate two buyer budgets at each time step $t$. The first buyer (high budget) has a budget uniformly randomly generated between $95\%$ and $100\%$ of the total utilities of all $10$ datasets listed in the market. The second buyer (low budget) has a budget uniformly randomly generated between $90\%$ and $95\%$ of the total utilities of all $10$ datasets listed in the market. At each time period, seller’s arriving orders are randomly shuffled. And they prices their own datasets sequentially.
Robustness Check. As discussed in Lemma 7, the threshold $t^{*}$ when fairshare becomes optimal for the buyer) increases as the discount factor $\delta$ decreases. To run a robustness check, for the high-budget buyer in Figure 4(a), setting $\delta=0.999,0.99,0.98$ yields $t^{*}=32,38,44$ respectively, which is consistent with our theoretical analysis. In below, we compare the change of buyer’s accumulative utilities over different values of $\delta$.
Figure 7: Buyer’s accumulative utilities over time periods for $\delta=0.980,0.990,0.999$.
Appendix F Additional Experiments and Figures
@@ F.1 Algorithms
1:Inputs: prices $\boldsymbol{p}$, previous optimal $\tilde{\boldsymbol{x}}^{k,N-1}$, feasible set $\mathcal{X}_{k,N-1}$, price $p_{j}$, budget $b_{k}$
2:Output: indicator $\mathbbm{1}_{\{B_{k},D_{j},p_{j}\}}$
3:Initialize $\mathbbm{1}_{\{B_{k},D_{j},p_{j}\}}\leftarrow 0$
4:$p_{j}\leftarrow 0$
5:for $\boldsymbol{x}\in\mathcal{X}_{k,N-1}$ do
6:     $\boldsymbol{x}^{\text{new}}\leftarrow\boldsymbol{x}+\boldsymbol{e}_{j}$
7:     if $g_{k,N}(\boldsymbol{x}^{\text{new}})>g_{k,N-1}(\tilde{\boldsymbol{x}}^{k,N-1})$ and $\boldsymbol{x}^{\top}\boldsymbol{p}+p_{j}\leq b_{k}$ then
8:         $\mathbbm{1}_{\{B_{k},D_{j},p_{j}\}}\leftarrow 1$
9:         break
10:     end if
11:end for
12:return $\mathbbm{1}_{\{B_{k},D_{j},p_{j}\}}$
@@
1:Inputs: Buyers $\{B_{k}\}_{k=1}^{M}$, Sellers $\{S_{j}\}_{j=1}^{N}$
2:Initialization: Buyers $\{B_{k}\}_{k=1}^{M}$ enter the market
3:for $j=1$ to $N$ do
4:     Seller $S_{j}$ enters with potential prices $\mathcal{P}_{j,M}$ for dataset $D_{j}$
5:     for all $p_{j}\in\mathcal{P}_{j,M}$ do
6:         for $k=1$ to $M$ do
7:              Buyer $B_{k}$ solves:
$\tilde{\boldsymbol{x}}^{k,j-1}=\arg\max_{\boldsymbol{x}\in\mathcal{X}_{k,j}}g_{k,j}(\boldsymbol{x})$
8:              to decide whether to purchase $D_{j}$ at price $p_{j}$triangleright See Eqn. 2
9:         end for
10:         Seller computes net profit $r(p_{j})$ assuming price $p_{j}$triangleright See Eqn. 3
11:     end for
12:     Seller selects:
$p_{j}^{*}=\arg\max_{p_{j}\in\mathcal{P}_{j,M}}r(p_{j})$
13:     and sets $p_{j}^{*}$ as the fixed price for $D_{j}$triangleright See Eqn. 4
14:end for
F.2 Data Valuation Oracle Experiments
Figure 8: Left, middle-left columns: Correlation analysis between oracle and $\text{Infl}_{\text{IP}}$ valuation. Right, middle-right columns: Performance analysis between between oracle and $\text{Infl}_{\text{IP}}$ valuation.
Influence-based methods, such as $\text{Infl}_{\text{IP}}$ approximate the influence of a sample $d$ on a $d^{\prime}$ for a model parameterized by $\theta$ by estimating the effects of training or “upweighting” (alternatively, removing/“downweighting”) on $d$ (see Appendix A). In past literature, $\text{Infl}_{\text{IP}}$ is validation through an “Oracle” (One-Step Training) score, which we denote as $\text{Oracle}(d,d^{\prime})=\mathcal{L}(d^{\prime};\theta)-\mathcal{L}(d^{\prime};\hat{\theta})$, where $\hat{\theta}=\theta-\eta\nabla\mathcal{L}(d;\theta)$ and $\eta$ is the learning rate [56, 62, 23].
To compare the difference between $\text{Infl}_{\text{IP}}$ valuation versus its oracle valuation, we conduct the same experiments described in Section 4.3. Figure 8 shows that $\text{Infl}_{\text{IP}}$ and Oracle have decent correlation in their agreement in their valuation of the sellers’ data, which supports findings in previous works [62]. We note that in the case where correlation is decent, such as in Figure 8(b), the final model performance between these methods is close, as seen in Figure 8(d). In the case where correlation is lower, such as in Figure 8(a), the final model performance between these methods initially have a gap, but become more similar as the amount of data purchased increases, as seen in Figure 8(c). This suggests that in practice, even in cases when the agreement between $\text{Infl}_{\text{IP}}$ and Oracle may not be very high, final model performance resulting from these two methods can still be similar.
F.3 Additional Experimental Results
Figure 9: Buyers’ model (top row: Pythia-1b, bottom row: Pythia-410m) performance and costs from their purchased data from math, medical, and physical reasoning data markets. Purchasing decisions were using the constant, random, BM25, $\text{Infl}_{\text{IP}}$ data valuation methods (see Section 4.3 for details).
Figure 10: Analysis of (1) buyer’s cumulative utilities with high-budget buyer (Figures 10(a), 10(e) and 10(i)) and low-budget buyer (Figures 10(b), 10(f) and 10(j)), and (2) sellers’ average cumulative profits (Figures 10(c), 10(g) and 10(k)) and number of sellers in the market (Figures 10(d), 10(h) and 10(l)) over time ($T=100$). Model: Pythia-1b, Pythia-410m, and Llama-3.2-Inst.-1b; Task: medqaQA. Experimental groups: (1) fairshare, (2) reduced, (3) random, and (4) current pricing.
Figure 11: Analysis of (1) buyer’s cumulative utilities with high-budget buyer (Figures 11(a), 11(e) and 11(i)) and low-budget buyer (Figures 11(b), 11(f) and 11(j)), and (2) sellers’ average cumulative profits (Figures 11(c), 11(g) and 11(k)) and number of sellers in the market (Figures 11(d), 11(h) and 11(l)) over time ($T=100$). Model: Pythia-1b, Pythia-410m, and Llama-3.2-Inst.-1b; Task: MathQA. Experimental groups: (1) fairshare, (2) reduced, (3) random, and (4) current pricing.
Figure 12: Analysis of (1) buyer’s cumulative utilities with high-budget buyer (Figures 12(a), 12(e) and 12(i)) and low-budget buyer (Figures 12(b), 12(f) and 12(j)), and (2) sellers’ average cumulative profits (Figures 12(c), 12(g) and 12(k)) and number of sellers in the market (Figures 12(d), 12(h) and 12(l)) over time ($T=100$). Model: Pythia-1b, Pythia-410m, and Llama-3.2-Inst.-1b; Task: PIQA. Experimental groups: (1) fairshare, (2) reduced, (3) random, and (4) current pricing.
F.4 Datasets